let/quiz?quiz-action=takeQuiz&quiz-probGuid-QNAPC0A80 1010000004091 d9d00 O0&ctx=morrism-00 3. Calculating risk and odds Aa Aa medical researcher investigates the relationship between the incidence of a particular disease and a known risk or for the disease in children under the age of 5. A random sample of 5,000 children under the age of 5 is studied. The data on their disease status (yes or no) and risk-factor presence (yes or no) is summarized in the following contingency table Has Risk Factor Yes No Total Yes 100 150 250 Has Disease No 1,150 3,600 4,750 Total 1,250 3,750 5,000 In this sample of children under the age of 5: The risk of having the disease is 0.050 The risk of having the disease if the child has the risk factor is The risk of having the disease if the child does not have the risk factor is A child with the risk factor is the risk factor. times more likely to have the disease than a child without In this sample of children under the age of 5: . The odds of having the disease are The odds of having the disease if the child has the risk factor are The odds of having the disease if the child does not have the risk factor are The odds of having the disease are times greater when the child has the risk factor than when the child does not have the risk factor.
Answer #1
  1. The risk of having the disease if the child has the risk factor = 100 / 5000 = 0.02
  2. The risk of having the disease if the child does not have the risk factor = 150 / 5000 = 0.03
  3. times more likely = (0.02 / 0.03) = 0.6667
  4. Odds of having the disease = 250 / 5000 = 1 / 20
  5. Odds of having the disease if the child has risk factor = 100 / 5000 = 1 / 50
  6. Odds of having the disease if the child does not have the risk factor = 150 / 5000 = 3 / 100
  7. times greater = (0.02 / 0.03) = 0.6667